# Whole-body Humanoid Robot Locomotion with Human Reference

### 28 Feb 2024 v1

@@ Abstract
Recently, humanoid robots have made significant advances in their ability to perform complex tasks due to the deployment of Reinforcement Learning (RL), however, the inherent complexity of humanoid robots, including the difficulty of planning complex reward functions and training entire complex systems, still poses a notable challenge. To conquer these challenges, after many iterations and in-depth investigations, we have meticulously developed a full-size humanoid robot, “Adam”, whose innovative structural design greatly improves the efficiency and effectiveness of the imitation learning process. In addition, we have developed a novel imitation learning framework based on an adversarial motion prior, which applies not only to Adam but also to humanoid robots in general. Using the framework, Adam can exhibit unprecedented human-like characteristics in locomotion tasks. Our experimental results demonstrate that the proposed framework enables Adam to achieve human-comparable performance in complex locomotion tasks, marking the first time that human locomotion data has been used for imitation learning in a full-size humanoid robot. For more video demonstrations, please visit our YouTube channel: https://www.youtube.com/watch?v=7hK2ySYBa1I
11footnotetext: The authors are with The Hong Kong University of Science and Technology (Guangzhou), China. $*{}^{*}$ are equal contributors, ${}^{\dagger}$ is the corresponding author {qzhang749, jsun444}@connect.hkust-gz.edu.cn, renjingxu@hkust-gz.edu.cn22footnotetext: The authors are with PNDbotics. {peter.cui, david.yan}@pndbotics.com33footnotetext: The author is with Human Centric AI Centre, Australia Artificial Intelligence Institute, University of Technology Sydney 2007 Ultimo Australia. yiqun.duan@student.uts.edu.au
@@ VI Conclusion
In this paper, we implement a whole-body imitation learning framework on the humanoid robot Adam, whose performance in complex gaits tasks is comparable to that of humans, and demonstrate human-like features such as ”heel-to-toe” transitions and straight-knee movements for the first time in humanoid robots. Our experiments have demonstrated the practicality and efficiency of our framework, highlighting its great potential for further research in humanoid robotics. Moreover, our framework supports the integration of perceptual modules, and in the future, we plan to incorporate more sensors to allow Adam to imitate human motions under integrated perceptual conditions.

### 01 Mar 2024 v2

@@ Abstract
Recently, humanoid robots have made significant advances in their ability to perform challenging tasks due to the deployment of Reinforcement Learning (RL), however, the inherent complexity of humanoid robots, including the difficulty of designing complicated reward functions and training entire sophisticated systems, still poses a notable challenge. To conquer these challenges, after many iterations and in-depth investigations, we have meticulously developed a full-size humanoid robot, “Adam”, whose innovative structural design greatly improves the efficiency and effectiveness of the imitation learning process. In addition, we have developed a novel imitation learning framework based on an adversarial motion prior, which applies not only to Adam but also to humanoid robots in general. Using the framework, Adam can exhibit unprecedented human-like characteristics in locomotion tasks. Our experimental results demonstrate that the proposed framework enables Adam to achieve human-comparable performance in complex locomotion tasks, marking the first time that human locomotion data has been used for imitation learning in a full-size humanoid robot. For more video demonstrations, please visit our YouTube channel: https://www.youtube.com/watch?v=7hK2ySYBa1I
11footnotetext: The authors are with The Hong Kong University of Science and Technology (Guangzhou), China. $*{}^{*}$ are equal contributors, ${}^{\dagger}$ is the corresponding author {qzhang749, jsun444}@connect.hkust-gz.edu.cn, renjingxu@hkust-gz.edu.cn22footnotetext: The authors are with PNDbotics. {peter.cui, david.yan, arthur.zhang}@pndbotics.com33footnotetext: The author is with Human Centric AI Centre, Australia Artificial Intelligence Institute, University of Technology Sydney 2007 Ultimo Australia. yiqun.duan@student.uts.edu.au
@@ VI Conclusion
In this paper, we implement a whole-body imitation learning framework on the humanoid robot Adam, whose performance in complex gaits tasks is comparable to that of humans, and demonstrate human-like features such as “heel-to-toe” transitions and straight-knee movements for the first time in humanoid robots. Our experiments have demonstrated the practicality and efficiency of our framework, highlighting its great potential for further research in humanoid robotics. Moreover, our framework supports the integration of perceptual modules, and in the future, we plan to incorporate more sensors to allow Adam to imitate human motions under integrated perceptual conditions.